If a and b are positive integers with a greatest common divisor of 3, and a/b=0.4, what is the product ab?
a 18 b 10 c 36 d 30 e 90

Answer by rajagopalan(174)

solution
let b=x
then a=0.4x
ab=x*0.4x=0.4x^2
now solving each a,b,c,d,e answers for 0.4x^2we get
1. 0.4x^2=18
x^2=18/0.4=45 this is not a perfect square. hence reject
2. 0.4x^2=10
x^2=10/0.4=25 this is a perfect square. but not divisible by 3. hence reject
3. 0.4x^2=36
x^2=36/0.4=90 this is not a perfect square. hence reject
4. 0.4x^2=30
x^2=30/0.4=75 this is not a perfect square. hence reject
5. 0.4x^2=90
x^2=90/0.4=225 this is a perfect square. x=15, so b=15, a=0.4x15=6
the answer is e